### v1

@@ 3.1 Analysis of collaboration diversity
Figure 2: Collaboration diversity over time based on (a) authors’ institutions and (b) authors’ departments. In both plots, the primary y-axis denotes the Shannon Entropy calculated using Equation 2, while the secondary y-axis indicates the count of papers. Panel (a) is based on 25,933 papers with complete institution information, while Panel (b) is based on 16,645 papers with complete department information. Each point in the panel represents the averaged entropy based on papers published in one year. The dotted gray line implies the date when ChatGPT was released by OpenAI.
3.2 Analysis of collaboration network
@@
We analyze the co-authorship networks in terms of structural cohesion, topology, community structures, and influential researchers with respect to various dimensions of centrality. The network metrics are presented in Table 1, and the network visuals are presented in Figure 3.
@@
Based on Table 1, the institution-based and department-based networks both have low density (0.0001) but high clustering coefficients (0.61 and 0.75, respectively), indicating that while direct collaborations are limited, existing ones form tight-knit clusters. The largest components comprise a notable portion of all network edges (77% for institutions network, and 70% for departmental affiliation network), meaning that most researchers are part of research communities that are reachable from each other.
@@
The relatively high average shortest path length (16.74 for institution-based, and 16.61 for department-based networks) suggests extensive reach within the networks, indicating that while direct connections between researchers might be limited, there are alternative paths that connect them. The primary difference is that the department-based network exhibits a stronger fit towards a power-law distribution ($\alpha$=2.32), suggesting that central departments, namely Computer Science, and Medicine, are major hubs that facilitate numerous collaborations. On the other hand, the institution-based network ($\alpha$=4.06) does not exhibit a clear hub-and-spokes structure, indicating a more evenly distributed pattern of collaborations across various institutions.
@@
Degree centrality results highlight the prominence of Computer Science, Medicine, and other Computer Science-related departments in LLM research. Furthermore, the high betweenness centrality of these same departments highlights their roles in facilitating collaboration between departments that would have otherwise been unconnected. As illustrated in Figure 3, the clusters of departments between co-authors not only demonstrate the dominance of Computer Science and Medicine and their related disciplines (the orange cluster and the green cluster for Computer Science and Medicine-related disciplines, respectively) in the network, but also shows that these two disciplines connect fields with little to no overlap. For instance, Medicine is in the shortest path between Engineering and Social Science in 196 instances. Similarly, Computer Science is in the shortest path between Medicine and Psychology and Neuroscience for 384 instances.
With respect to key institutions (the “Degree Centrality” and “Betweenness Centrality” rows in Table 1), Stanford University is consistently the highest in terms of degree and betweenness centrality, indicating their active involvement as both collaborators and facilitators of collaboration between institutions in LLM research. While Harvard University and University College London are not as central in terms of direct collaborations with other universities, their high betweenness indicates they actively connect institutions with each other, as well as with industrial tech companies: for instance, University College London connects Google (United States) and University of Zurich (162 instances); and Harvard University connects Allen Institute and University of Washington (210 instances).
Closeness centrality results show how researchers from departments and institutions are, on average, reachable to other researchers. Top 5 departments with the highest closeness centrality are in language-related disciplines, such as Computational Language Modeling, German Language, and Literature & Education, which shows the extensive application of LLMs in domains where language analysis or content generation is needed. Top 5 institutions in terms of closeness centrality are primarily universities from countries outside the U.S. (see Table 1). This finding shows that LLM-related research is a global effort, with international institutions acting as connectors across geographical boundaries.
Eigenvector centrality results show more consistency with degree centrality and betweenness centrality results than with closeness centrality, highlighting researchers who are not only well-connected but also connected to other influential researchers. Researchers with the highest eigenvector centrality scores are also from Medicine, Computer Science, and their related fields. Biomedical Informatics has the highest eigenvector centrality, as it is directly connected to Medicine and Computer Science, the two most influential departments in our dataset. The eigenvector centrality analysis for institutions underscores the critical role of medical research institutes, such as the University of Colorado Anschutz Medical Campus, European Bioinformatics Institute, and Lawrence Berkeley National Laboratory, in facilitating collaborations that significantly impact the LLM research community.
@@
Figure 3: Co-authorship networks based on authors’ (a) institution and (b) department affiliations. Each node represents a researcher’s institution or department, and each edge represents a co-authorship between pairs of researchers from the respective institutions or departments. Node colors represent the clusters to which the nodes belong, determined based on the Louvain modularity. Node labels represent the top 20% of nodes ranked by degree centrality. Edge thickness represents the frequency of co-authorship between the connected researchers.
@@
Table 1: Network metrics based on authors’ institutions and departments
@@
No. Nodes 8391 6604
No. Edges 35080 16811
Density 0.001 0.0008
Avg. Shortest Path 16.74 16.61
Clustering Coeff. 0.61 0.75
No. Components 1569 1678
Largest Component 6485 4616
Power-law Exponent 4.06 2.32
No. Comm.(Louvain) 1613 1730
No. Comm. (CNM) 1684 1775
Degree Centrality Stanford University (343) Computer Science (835)
(top 5) University of Washington (287) Medicine (518)
Peking University (262) Computer Science & Eng. (422)
Tsinghua University (256) Artificial Intelligence (289)
University of Waterloo (227) Engineering (226)
Betweenness University College London (0.028) Computer Science (0.089)
Centrality Harvard University (0.026) Computer Science & Eng. (0.038)
(top 5) Stanford University (0.021) Medicine (0.035),
University of Waterloo (0.021) Psychology (0.034)
University of Washington (0.021) Engineering (0.034)
Closeness Centrality University of Craiova (0.5) Computational Lang. Modeling (0.50)
(top 5) Constantin Brâncuşi Univ. (0.5) German Language (0.50)
Nutrition Center of Philippines (0.5) Information & Architecture (0.50)
West Visayas State University (0.5) Literature & Education (0.50)
Yalova University (0.5) Vet. Public Health & Epidem. (0.50)
Eigenvector Univ. of Colorado Anschutz (0.61) Biomedical Informatics (0.45)
Centrality Lawrence Berkeley Natl. Lab. (0.54) Computer Science (0.40)
(top 5) European Bioinfo. Institute (0.37) Medicine (0.38)
Critical Path Institute (0.31) Envir. Genomics & Sys. Biology (0.30)
University of Illinois System (0.11) Bioinformatics (0.26)
@@ 4.1 Key findings and implications
Collaboration diversity results, as well as co-authorship structures, reveal the notable and complex effects that LLMs have on fostering interdisciplinary and cross-institutional collaborations. The integration of LLMs into research areas has led to a potential shift in collaboration patterns. Below, we summarize several key findings and implications.
@@
First, there has been an overall significant increase in the number of publications on LLM-related topics since the advent of ChatGPT in 2022. This surge extends beyond Computer Science and artificial intelligence to disciplines such as Medicine, Social Science, and Engineering. Additionally, the overall entropy, which measures the diversity of collaboration, has increased. This observation implies that collaboration diversity has broadened across these fields.
The impact on specific fields varies. Computer Science sees a stable increase in entropy, suggesting that computer scientists continue to seek collaboration with researchers from other disciplines. This trend aligns with our observation that researchers are increasingly exploring topics in AI for social good or AI for science [27]. Health-related fields, including Medicine, Neuroscience, Health Professions, and Biochemistry, Genetics, & Molecular Biology, exhibit higher entropy compared to Computer Science, Social Sciences, and Engineering. However, Medicine is the only area showing a significant decrease in entropy. One possible explanation is that the advancement of LLMs, particularly with their simple deployment interfaces, could enable researchers from disciplines like Medicine to more effectively use these AI models for domain-specific issues compared to the past. This is already being seen in applications such as drug discovery and repurposing [28].
@@
Therefore, while LLMs might be associated with broader collaborations across disciplines and institutions, they may also encourage specialized collaborations within certain fields, such as Medicine, where domain-specific knowledge is often needed. In particular, medical-related research may have specific data privacy requirements, such as HIPAA regulations for electronic health records [29], as well as domain-specific expertise, such as in pathology [30] and clinical diagnosis [31], to evaluate whether LLM-predicted outputs are correlated with improved patient outcomes. Unlike previous AI models, researchers in Medicine can now use LLMs to address domain-specific issues more effectively without always seeking external collaborations.
These findings suggest that interdisciplinary collaboration may foster in contexts where LLMs may be used for analyzing and generating data that could be applicable to multiple disciplines, as opposed to health-related fields like Medicine where the data is typically specific to that field. Researchers in these disciplines can now use LLMs to address domain-specific issues more effectively without always seeking external collaborations. Additionally, creating clear guidelines and support for managing data privacy, ethical considerations, and domain-specific adaptations of LLM technologies can be crucial for leveraging their full potential in sensitive fields like Medicine.
@@
Structural analysis of co-authorship connections shows that despite differences in collaboration diversity after the advent of ChatGPT, Computer Science and Medicine remain the most represented disciplines in the network. In particular, researchers from these two disciplines (via departmental affiliation) have the highest degree and betweenness centrality scores, indicating their roles as both active researchers and facilitators of cross-field collaborations. This finding is especially important with regard to the field of Computer Science because, despite their own papers having less disciplines represented (less entropy compared to other disciplines), they are structurally necessary to connect other disciplines together. For instance, Medicine bridges Computer Science and Engineering and Information Technology (334 times), Engineering and Social Science (196 times), and Neurosurgery and Intelligent Technology and Engineering (144 times). Relatedly, the influence of Medicine is also reflected in the cross-institutional analysis, where medical institutions and associated national laboratories (e.g. University of Colorado Anschutz Medical Campus, European Bioinformatics Institute, Jackson Laboratory) are the most influential in terms of eigenvector centrality, highlighting their central roles in connecting with other influential institutions and facilitating extensive cross-institutional collaborations.
@@
Overall, the study of author collaboration and network analysis provides valuable insights into the dynamics and patterns of interdisciplinary research in LLMs, highlighting how knowledge and expertise are exchanged across various fields. Our analysis reveals that the diversity of academic collaborations increases overall, but varies across disciplines. Fields with methodological expertise and domain generality like Computer Science exhibit an increasing collaboration diversity as they might have sought for applications with other different fields, such as Psychology and Social Sciences. In contrast, disciplines like Medicine show a decreased diversity in collaboration after the release of ChatGPT due to the necessity of domain-specific knowledge in addition to methodological expertise and the more effective application of these AI models compared to the past. Extensive cross-institutional and cross-country collaborations are also facilitated by two dominant fields, Computer Science and Medicine. Leading research institutions and associated national labs in the U.S., U.K., and China are actively conducting research using LLMs and fostering interdisciplinary partnerships while sharing expertise and resources.

### v2

@@ 3.1 Analysis and comparison of collaboration diversity
Figure 2: Collaboration diversity over time based on (a) authors’ institutions and (b) authors’ departments. Panel (a) is based on 30,075 papers with complete institution information, while Panel (b) is based on 22,453 papers with complete department information. In both plots, the x-axis denotes the date by year, while the y-axis denotes the Shannon Entropy calculated using Equation 2. Each point represents the monthly averaged entropy based on papers published in one quarter. The line represents the trend using a non-parametric regression technique called locally weighted scatterplot smoothing (LOWESS). Each point in the panel represents the averaged entropy based on papers published in one year. The dotted gray line implies the date when ChatGPT was released by OpenAI.
3. How did the release of ChatGPT influence collaboration dynamics? Our next question explores the impact of ChatGPT’s release. As presented in Figure 2, the variance of entropy is much wider before ChatGPT’s release, possibly due to fewer papers on the topic of LLMs and more diverse collaborations. The averaged entropy is high for some fields like Engineering. One possible explanation for this pattern is that researchers from other fields sought partnerships with Computer Science experts for such research topics. In addition, the observed changes in entropy suggest dynamics across different fields. Papers in the field of Computer Science display a stable increasing trend. Medicine, Social Science, Health Professions, and Decision Science also exhibit a notable increase in institution entropy. Other fields display either minor increases or negligible changes. These findings potentially imply the varying impacts of ChatGPT’s introduction on interdisciplinary collaborations across different academic fields.
We conduct a Wilcoxon rank-sum test to compare the entropy before and after the release of ChatGPT, with detailed results presented in Appendix B (see Figure 7). Our analysis reveals statistically significant changes in entropy across several fields. Computer Science, Social Science, Psychology, and Biochemistry, Genetics & Molecular Biology show a statistically significant increase in both institution and department entropies, while Medicine and Neuroscience display a slight, but not statistically significant, decrease. The other fields do not show any significant change in entropy. In particular, we observe that the mean of the entropy distribution for most fields remains at 0, with the first quartile consistently at 0. This suggests that many researchers primarily collaborate with colleagues from their own institutions or departments on LLM-related topics.
We conduct additional Wilcoxon rank-sum tests to compare entropy across different fields, with results presented as heatmaps in Figure 8. Several consistent patterns emerge when examining both institution and department results before and after ChatGPT’s release. Certain fields like Computer Science, Medicine, and Biochemistry, Genetics & Molecular Biology exhibit more significant differences compared to other fields, as evidenced by darker color cells in the heatmaps. Following ChatGPT’s release, there is also an increase in significant differences across fields, illustrated by a higher prevalence of darker orange cells in the right heatmap in Figure 8, compared to their counterparts in the left heatmap. Another interesting observation is that Computer Science consistently demonstrates significantly lower entropy than Medicine, Biochemistry, Genetics & Molecular Biology, Health Professions, and Neuroscience.
We further perform difference-in-difference (DiD) analysis to examine the impact of ChatGPT. Figure 3 displays results of $\beta_{3}$ for the top 12 fields. As shown, the release of ChatGPT consistently leads to a statistically significant increase in entropy in fields such as Computer Science and Social Science, and a statistically significant decrease in entropy in fields such as Economics, Econometrics, and Finance. It also leads to a statistically significant increase in entropy in Health Professions and Decision Sciences based on institutions, and a statistically significant increase in entropy in Biochemistry, Genetics, and Molecular Biology based on departments. Other fields, however, do not show consistent and significant causal impacts from ChatGPT. The strongest positive impacts are mainly observed in health-related fields, such as Health Professions and Biochemistry, Genetics, and Molecular Biology.
Figure 3: The impact of ChatGPT on the entropy of LLM group as compared to (a) ML and (b) Non-LLM/ML paper groups. Coefficients and Std. Error of DiD ($\beta_{3}$) by field. When the confidence interval does NOT cross zero (the dashed line), the effect is considered statistically significant, implying that we can be 95% confident that ChatGPT has a real impact that isn’t due to random chance.
3.2 Analysis and comparison of collaboration networks
@@
We analyze the co-authorship networks regarding structural cohesion, topology, community structures, and influential researchers, focusing on dimensions of centrality compared to the control groups. The network metrics appear in Table 1, and the network visuals appear in Figure 4-Figure 5. Our analysis in this section aims to explore the folllowing three questions in terms of collaboration structure.
@@
1. How do the co-authorship networks for LLM papers compare to those in ML and non-LLM/ML research? Based on Table 1, both the institution-based and department-based networks show low density (0.001) and high clustering coefficients (0.60 and 0.79, respectively), but existing ones form tight-knit clusters. The largest components make up a significant part of all network edges (78% for the institutions network and 76% for the departmental affiliation network). This suggests that most researchers belong to research communities that can reach one another. The control groups’ networks resemble the LLM network, exhibiting low density (range: 0.0003-0.001) and high clustering coefficients (range: 0.59-0.81).
@@
The relatively high average shortest path length (14.37 for institution-based, and 18.57 for department-based networks) suggests extensive reach within the networks, indicating that while direct connections between researchers might be limited, there are alternative paths that connect them. The LLM-based network, similar to the control group’s networks, also exhibits a stronger fit towards a power-law distribution ($\alpha$ range: 2.24-3.75), suggesting that central departments, namely Computer Science, and Medicine, are hubs that facilitate numerous collaborations across all networks, regardless of whether the collaborations are on topics that are LLM-related, ML-related, or non-LLM-related. Table 5 reveals that the top central nodes across domains primarily consist of computer science, medicine, and their related subdisciplines.
@@
2. How do the top institutions and departments in LLM papers differ from those in ML and non-LLM/ML research? Degree centrality results highlight the prominence of Computer Science, Medicine, and other Computer Science-related departments in LLM research. Furthermore, the high betweenness centrality of these same departments highlights their roles in facilitating collaboration between departments that would have otherwise been unconnected. As shown in Figure 5, the clusters of departments between co-authors not only demonstrate the dominance of Computer Science and Medicine and their related disciplines (the orange cluster and the green cluster for Computer Science and Medicine-related disciplines, respectively) in the network, but also shows that these two disciplines connect fields with little to no overlap. For instance, Medicine is in the shortest path between Engineering and Social Science in 196 instances. Similarly, Computer Science is in the shortest path between Medicine and Psychology and Neuroscience for 384 instances.
The betweenness centrality results reveal notable differences, with international institutions, such as Instituto de Ciencias de La Construcción Eduardo Torroja and Heilongjiang Academy of Sciences, serving as key bridges in the LLM network. These institutions connect otherwise disconnected research communities, particularly linking major hubs like Chinese Academy of Sciences, Tsinghua University, and University of Amsterdam with smaller or regional institutions. This contrasts with the ML and non-LLM/ML networks, where well-established institutions, primarily from the U.S. and U.K., dominate. At the department level, the LLM network highlights more specialized areas, such as Dynamics, Logics and Inference for Biological Systems and Data and Knowledge Management, while traditional departments like Computer Science and Medicine continue to hold central positions in the ML and non-LLM/ML networks, reflecting their long-standing influence between communities of research.
Closeness centrality results show how researchers from departments and institutions are, on average, reachable to other researchers. In the LLM network, prominent colleges across U.S. (Harvard University), the U.K. (UCL), and China (Peking University) are on average, closest to all other institutions in the network. On the other hand, the ML and non-LLM/ML networks are more fragmented, with smaller or regional institutions such as Trilogi University and Universitas Bina Darma with the highest closeness scores. At the department level, Computer Science and Medicine are the most central in the LLM network, while in the ML and non-LLM/ML networks, Computer Science-related departments dominate, with no Medicine-related departments appearing as central in terms of closeness centrality.
Eigenvector centrality results show more consistency with degree centrality and betweenness centrality results than with closeness centrality, highlighting researchers who are not only well-connected but also connected to other influential researchers. Institutions with the highest eigenvector centrality scores are also those with high degree and betweenness centrality, namely Stanford University and Tsinghua University.
@@
Figure 4: Co-authorship networks based on authors’ institutional affiliations. Each node represents a researcher’s institution, and each edge represents a co-authorship between pairs of researchers from the respective institutions or departments. Node colors represent the clusters to which the nodes belong, determined based on the Louvain modularity. Node labels represent the top 20% of nodes ranked by degree centrality. Edge thickness represents the frequency of co-authorship between the connected researchers.
Figure 5: Co-authorship networks based on authors’ department affiliations. Each node represents a researcher’s department, and each edge represents a co-authorship between pairs of researchers from the respective institutions or departments. Node colors represent the clusters to which the nodes belong, determined based on the Louvain modularity. Node labels represent the top 20% of nodes ranked by degree centrality. Edge thickness represents the frequency of co-authorship between the connected researchers.
@@
Table 1: Comparison of network metrics based on authors’ institutions and departments
@@
LLM ML Non-LLM/ML LLM ML Non-LLM/ML
No. Nodes 9320 16391 19010 10131 24707 26492
No. Edges 40971 119960 104649 33314 121854 94901
Density 0.0009 0.0009 0.0006 0.0006 0.0004 0.0003
Avg. Shortest Path 14.37 13.4 15.36 18.57 16.47 22.79
Clustering Coefficient 0.6 0.59 0.6 0.79 0.8 0.81
No. Components 1729 1481 3113 1963 2786 4686
Largest Component 7266 14699 15426 7679 21261 20649
Power-law Exponent 3.75 2.29 2.54 2.29 2.24 2.31
No. Communities (Louvain) 1774 1508 3150 2010 2833 4729
No. Communities (CNM) 1856 1707 3308 2091 3126 5066
3. How similar (or different) are LLM networks to ML and non-LLM/ML networks in node, edge, and structural overlap? At the institutional level, the LLM and ML networks share notable node overlaps, such as Stanford University, Tsinghua University, and University College London, which appear as central nodes in both networks. However, the edge Jaccard similarity remains low (0.025), showing that while these institutions are present in both networks, their patterns of collaboration differ significantly. In contrast, the LLM network shares fewer nodes with the non-LLM/ML network, as seen in the lower node Jaccard similarity (0.293), with some overlap in institutions like the Chinese Academy of Sciences and Peking University. At the overall network level, the Netsimile distance indicates that the LLM network is more similar to the non-LLM/ML network (3.203) compared to the ML network (6.275), suggesting that LLM research draws from a broader institutional base rather than mirroring the more concentrated ML collaborations.
At the department level, node and edge Jaccard values are consistently lower than those at the institutional level, with LLM to ML showing slightly higher overlap (0.141 for nodes, 0.03 for edges) compared to LLM to non-LLM/ML (0.113 for nodes, 0.02 for edges). This indicates that while Computer Science is central across all networks, the LLM network includes more specialized departments like Artificial Intelligence, Data Science, and Information Science, which do not appear prominently in the ML and non-LLM/ML networks. In contrast, the non-LLM/ML network contains nodes representing traditional fields such as Physics and Biological Sciences, while the ML network is concentrated around Computer Science, Engineering, and Electrical Engineering. The Netsimile distance further highlights these differences, showing greater structural similarity between LLM and non-LLM/ML networks (2.264) than between LLM and ML networks (6.35). This suggests that ML-driven collaborations are more tightly clustered around established technical fields, whereas the LLM network, like the non-LLM/ML network, it contains clusters from a broader range of domains, spanning computational sciences to environmental sciences and business.
Table 2: Comparison of node Jaccard similarity, edge Jaccard similarity, and Netsimile distance
Comparison Node Jaccard Similarity Edge Jaccard Similarity Netsimile Distance
Institutions
LLM to ML 0.332 0.025 6.275
LLM to non-LLM/ML 0.293 0.019 3.203
Departments
LLM to ML 0.141 0.03 6.35
LLM to non-LLM/ML 0.113 0.02 2.264
@@ 4.1 Key findings and implications
Collaboration diversity results, as well as co-authorship structures, reveal that LLMs have notable and complex effects on interdisciplinary and cross-institutional collaborations. Integrating LLMs into research areas has the potential to shift collaboration patterns. Below, we summarize several key findings and implications.
@@
First, scientific collaboration generally increases in frequency, but the specific patterns vary by discipline. Since the advent of ChatGPT in 2022, the number of publications on LLM-related topics has increased significantly. This surge extends beyond Computer Science to fields such as Medicine, Social Science, and Engineering. Additionally, the overall entropy, which measures collaboration diversity, has increased. This observation implies that collaboration diversity has broadened across fields that employ LLMs in research.
However, the entropy pattern on specific fields varies. Computer Science sees a stable increase in entropy, suggesting that computer scientists continue to seek collaboration with researchers from other disciplines. This trend aligns with the observation that researchers explore topics such as AI for social good and AI for science [44]. Social Science also exhibits a significant rise in both institution entropy and department entropy following the release of ChatGPT. This likely reflects the growing application of computational methods and AI tools among social scientists to conduct quantitative analyses [45]. Health-related fields, including Medicine, Neuroscience, Health Professions, and Biochemistry, Genetics, & Molecular Biology, exhibit higher entropy compared to Computer Science, Social Sciences, and Engineering. Among these health-related fields, Psychology, Biochemistry, Genetics, & Molecular Biology, and Health Professions have experienced particularly notable increases in institution entropy. Our observation aligns with prior research showing that researchers in health are enthusiastic about using LLMs to facilitate documentation of patient reports, improve diagnosis accuracy, and assist in clinical care [46]. It is worth noting that Computer Science exhibits lower entropy overall compared to health-related fields. One possible explanation is that health-related articles often have longer author lists. Additionally, in the early stages of LLM research, researchers from various disciplines sought collaborations with Computer Science researchers. This focus aligned with the fact that LLMs were initially concentrated within the domain of Computer Science, as it was a highly specialized topic during that period.
Second, scientific collaboration within the LLM paper group indicates potential for increased interdisciplinary collaboration. This trend is particularly evident in fields such as Computer Science, Medicine, Social Science, and Health Professions. In general, AI-related studies exhibit higher levels of collaboration, with ML paper groups consistently positioned above Non-LLM/ML paper groups. However, the entropy of the LLM paper group starts at a much lower level, indicating a higher entry barrier in earlier stages. After the release of ChatGPT, collaboration across most fields demonstrates an upward trend, and the gap between the LLM paper group and the ML paper group narrows.
A plausible explanation is that Computer Science researchers increasingly seek partnerships with domain experts or are sought by their colleagues focusing on interdisciplinary areas such as AI for scientific applications or AI for social good. In other fields, the advancements in LLMs may have lowered entry barriers or introduced new opportunities in domain-specific areas. For instance, domain-specific studies in pathology [47] and clinical diagnosis [46], which evaluate whether LLM-predicted outputs correlate with improved patient outcomes, likely contribute to the growth of interdisciplinary collaboration.
Third, the advancement of LLMs has a significant impact on scientific collaboration. Our analysis indicates that these advancements like the release of ChatGPT have potentially lowered the entry barriers and broadened the applications of LLMs across various fields. Consequently, most disciplines have shown a substantial increase in the number of publications. Furthermore, many fields, including Computer Science and Social Science, have witnessed a consistent increase in the diversity of collaborations.
@@
For other fields like Medicine and Neuroscience, inter-institutional collaboration has grown, but intra-departmental collaboration has not increased similarly. While LLMs have facilitated broader collaborations across institutions, they may also promote specialized collaborations within certain domains where expert knowledge is critical. For example, medical research often involves stringent data privacy requirements, such as HIPAA regulations for electronic health records [48]. Unlike earlier AI models, researchers in Medicine can now leverage LLMs to address domain-specific challenges more effectively, often without requiring external collaborations from other departments.
@@
Fourth, structural analysis of co-authorship connections shows that Computer Science and Medicine remain the most represented disciplines in the network. In particular, researchers from these two disciplines have the highest degree and betweenness centrality scores, indicating their roles as both active researchers and facilitators of cross-field collaborations. This finding holds special importance for the field of Computer Science because, despite their papers featuring fewer disciplines (indicating less entropy compared to others), they play a crucial role in connecting various disciplines. For instance, Medicine bridges Computer Science and Engineering and Information Technology, Engineering and Social Science, and Neurosurgery and Intelligent Technology and Engineering. Medicine’s influence is also reflected in the cross-institutional analysis, where medical institutions and associated national laboratories (e.g. University of Colorado Anschutz Medical Campus, European Bioinformatics Institute, Jackson Laboratory) are the most influential in terms of eigenvector centrality, highlighting their central roles in connecting other influential institutions and facilitating extensive cross-institutional collaborations.
Last, structural analysis reveals that the advancement of LLMs has the potential to foster interdisciplinary collaboration. The LLM collaboration network preserves network properties similar to established research fields while showcasing a unique democratization of research participation. International institutions and specialized departments play prominent roles as key bridges within the network, marking a significant shift from the traditional ML research paradigm, which historically centered around leading U.S. and U.K. institutions. The lower Jaccard similarity values between LLM and ML networks suggest that ChatGPT has sparked potential cross-disciplinary collaborations, reaching beyond traditional ML research boundaries. Additionally, the high clustering coefficients, coupled with the network’s extensive reach, indicate that researchers still collaborate within their immediate communities. However, ChatGPT’s versatility has enabled researchers to forged connections between previously disconnected domains. This shift has fostered a more distributed and interdisciplinary research ecosystem. These findings highlight how ChatGPT has not only democratized access to LLM research but also actively facilitated the growth of interdisciplinary collaborations across diverse fields.
@@
Overall, the study of author collaboration and network analysis provides valuable insights into the dynamics and patterns of interdisciplinary research in LLMs, highlighting how knowledge and expertise are exchanged across various fields. Our analysis reveals that the diversity of academic collaborations increases overall, but varies across disciplines. Fields with methodological expertise and domain generality like Computer Science exhibit an increasing collaboration diversity. In contrast, disciplines like Medicine exhibit decreased collaboration diversity after the release of ChatGPT. This decline likely results because LLMs lower the demands for external technical and/or methodological expertise. Two dominant fields, Computer Science and Medicine, facilitate extensive cross-institutional and cross-country collaborations. Leading research institutions and associated national labs in the U.S., U.K., and China actively conduct research using LLMs, fostering interdisciplinary partnerships while sharing expertise and resources.